what is the answer to
14-2z=16
2n-1/3=10

13-5y=8

Answer by checkley71(8403)

14-2z=16 or -2z=16-14 or 2z=-16+14 or 2z=-2 or z=-1.
2n-1/3=10 or 2n=10+1/3 or 2n=10 1/3 or n=10 1/3 divided by 2 or n=5 1/6.
13-5y=8 or -5y=8-13 or 5y=-8+13 or 5y=5 or dividing both sides by 5 yields y=1.